OpenStudy (anonymous):

shirts increased from 20 to 26 what was rise in percentage

OpenStudy (luigi0210):

Hint: 10% =2

OpenStudy (anonymous):

percentage rise = 6/20 = 30/100 = ?

OpenStudy (primeralph):